OpenStudy (anonymous):

sketch the graph bounded by the graphs of the given equation and find the area of that region y=x^3+1 y=x-1 x=-1 x=1 this is area between curves

OpenStudy (amistre64):

for starters; add +2 to it all to get it above the xaxis;

OpenStudy (anonymous):

http://lj010.k12.sd.us/graph%20x%5E3.gif thats the x^3 and you have to add +1 to the y

OpenStudy (anonymous):

the x-1 is a line and you have to subtract 1

OpenStudy (amistre64):

or you might be able to just get away with shifting it all tothe right by 1 :)

OpenStudy (anonymous):

huh :S?

OpenStudy (anonymous):

i know how the graphs look i just dont know where they touch and...which is upper and which is lower

OpenStudy (anonymous):

for the intersection points just put (x-1) in y

OpenStudy (anonymous):

It is possibly an error in writing down the question or an error by your instructor: There is no defined error. It goes to infinity.

OpenStudy (anonymous):

no defined area, I meant to say.

OpenStudy (anonymous):

is the x=-1 and x=1 points of intersection?

OpenStudy (anonymous):

OK, I got it the x's are considered boundaries.

OpenStudy (anonymous):

oh wow thanks amistre :)

OpenStudy (anonymous):

i drew the same thing...

OpenStudy (anonymous):

in the boundet area of x=-1 to x =1 there are no intersection points

OpenStudy (amistre64):

i dont think the part beneath the xaxis hurts you; but you might wanna shift all this to the right by 1

OpenStudy (anonymous):

Don't bother shifting Amistre, it works on in the wash.

OpenStudy (amistre64):

y=(x-1)^3+1 y=(x-1)-1 y=x^3 -3x^2 +3x -2 y=x-2

OpenStudy (amistre64):

:) i hope so; i can never remember what the outcomes are with left and right of the y axis

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}[(x ^{3}+1)-(x -1)]\]

OpenStudy (anonymous):

how do you know which is upper and lower?

OpenStudy (anonymous):

You can graph it. If you cant use a calculator, you should know the x^3 starts at 1 and goes up and the x starts at -1. So x^3 is on top.

OpenStudy (anonymous):

what on top okay...

OpenStudy (amistre64):

y=x^3 -3x^2 +3x y=x-2 {S} (x^3 -3x^2 +3x - (x-2) dx ; [0,2] {S} (x^3 -3x^2 +2x +2) dx ; [0,2] x^4/4 -x^3 +x^2 +2x ; at x=0 is 0 (2)^4/4 -(2)^3 +(2)^2 +2(2) (2)^4/4 -8 +4 +4 2.2.2.2 ------ = 4 if i did it right lol 2.2

OpenStudy (anonymous):

i got 4 too

OpenStudy (amistre64):

andit 4 without the shift ;)

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}(x^3+1)-(x-1) dx\]

OpenStudy (amistre64):

if i move my graph like puzzle pieces I can make 4 boxes too lol

OpenStudy (anonymous):

i did it using that way though

OpenStudy (amistre64):

purp; now do it with the functions reversed {S} (x-1) - (x^3+1) dx {S} x-1 -x^3-1 dx {S} x -x^3-2 dx x^2/2 -x^4/4 -2x

OpenStudy (anonymous):

lol thanks so much amistre, chaguanas and nistal! :)

OpenStudy (amistre64):

it should be the same answer only negative

OpenStudy (anonymous):

oh ok cool...

OpenStudy (amistre64):

so dont worry so much about which functions on top or bottom just take the |results| :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok thats if we dont have a question sepcifically on area under below x axis or area above x-axis lol